MATLAB (Matrix Laboratory) —

The computational environment we should be universally expecting to learn and use.

Basic Matlab:

(a) The Mother of all calculators — where everything can be done with matrices

(b) Programming language — very powerful and much easier than, for example, C or Fortran (no array dimensioning)

(c) Graphical User Interface Development Environment (GUIDE) — do your own data collection GUIs or demonstrations

(d) Provides great graphics — plotting, visualization, animation, image processing

(e) Application Program Interface (API) — for calling C, Fortran, JAVA routines within the Matlab environment Toolboxes (collection of m-files to do analyses within some specific area, and where typically the source code in the m-files is available for modification)

Commercial (though the Mathworks or thirdparties) Toolboxes —

Statistics; Optimization; Image Processing; Neural Networks; Curve Fitting; Splines; Wavelets; Symbolic Math; Bioinformatics; Genetic Algorithm and Direct Search; Mapping (Geographical);

Compiler (m-files direct to C or C++ code)

Free (open sources) Toolboxes —

www.mathworks.com/matlabcentral/ — 1000 m-files contributed by users

www.mathtools.net/MATLAB/toolboxes.html

Auditory Perception; ILAB (eye movement analyses); Brainstorm (MEG/EEG); Chemometrics; N-way Toolbox; Smoothing; Econometrics; Spatial Statistics; NURBS; ICA (EEG/ERP – independent component analysis); LYNGBY (fMRI analyses); NIH CORTEX; NMRLAB; Psychophysics; Speech Processing; EMEGS (for EEG and MEG)

The Structural Representation of Proximity Matrices With MATLAB — This is the Combinatorial Data Analysis (CDA) Toolbox I’m working on and is available at (my office on the fourth floor): http://cda.psych.uiuc.edu/srpm mfiles — all the m-files I’ve written for the toolbox all the *.dat files for the data analysis examples plus the (draft) manual — cda toolbox manual rev1.pdf As two examples of how MATLAB has become the medium for sharing work, check out — http://cda.psych.uiuc.edu/lee files http://cda.psych.uiuc.edu/lee site http://cda.psych.uiuc.edu/Res Also, I have made all the relevant Matlab pdfs available at: http://cda.psych.uiuc.edu/matlab pdf